A toroid is made of 2 000 turns of wire of radius 2.00 cm formed into a donut shape of inner radius 10.0 cm and outer radius 14.0 cm. When a 10.0 -A current is present in the toroid , determine the magnetic field at a distance of 11.0 cm from the centre of the toroid.
1
Expert's answer
2020-03-31T09:24:33-0400
"B=\frac{\mu_0NI}{2\pi r}"

"B=\frac{(4\pi \cdot 10^{-7})(2000)(10)}{2\pi (0.11)}=0.0364\ T"